OpenStudy (anonymous):

A jar contains jelly beans that are red, yellow, blue, and purple. Several people reach in and grab handfuls of jelly beans, then collectively tally the number of each color: 15 red, 20 yellow, 12 blue, and 8 purple. If someone pulls out 22 jelly beans, about how many would you expect to be yellow? A. 20 B. 4 C. 8 D. 13

OpenStudy (anonymous):

@Thesmarterone

OpenStudy (anonymous):

20/55 = x/22 440 = 55x 8=x

OpenStudy (anonymous):

so...? answer is what?

OpenStudy (anonymous):

hm

OpenStudy (anonymous):

sory my computer messed up

OpenStudy (anonymous):

c